a charged rod of length l and having charge q revolves about its one end with frequency ω.its magnetic moment is ____(ans:qωl2/6)

Grade:12

1 Answers

ROSHAN MUJEEB
3 years ago
M/L=q/2m. M=qwl^2/6(by putting value of m,q and L)